Two years ago a father was four times as old as his son. Three years from now the father will be only three times as old as the son. How old is each now?
Answer by ankor@dixie-net.com(22740)

f = father's present age
s = son's present age
:
write an equation for each statement, simplify
Two years ago a father was four times as old as his son.
f - 2 = 4(s - 2)
f - 2 = 4s - 8
f = 4s - 8 + 2
f = 4s - 6
Three years from now the father will be only three times as old as the son.
f + 3 = 3(s + 3)
f + 3 = 3s + 9
f = 3s + 9 - 3
f = 3s + 6
:
Replace f with (4s-6) from the 1st equation
4s - 6 = 3s + 6
4s - 3s = 6 + 6
s = 12 yrs is the son's age
:
Use the equation; f = 3s + 6 to find f
f = 3(12) + 6
f = 42 yrs is father's age
:
How old is each now? 42 & 12
:
:
You can check to see if this is true in both statements